OpenStudy (anonymous):

For the following system, use the second equation to make a substitution for y in the first equation. 2x + y = 5 y + 3 = 2x

OpenStudy (dumbcow):

y = 2x-3 -> 2x + (2x-3) = 5 4x -3 = 5 4x = 8 x = 2 y = 2(2) -3 y = 4-3 y= 1